OpenStudy (anonymous):

31.5 grams of an unknown substance is heated to 102.4 degrees Celsius and then placed into a calorimeter containing 103.5 grams of water at 24.5 degrees Celsius. If the final temperature reached in the calorimeter is 32.5 degrees Celsius, what is the specific heat of the unknown substance? Show OR explain the work needed to solve this problem, and remember that the specific heat capacity of water is 4.18 J/(degrees C * g).

OpenStudy (anonymous):

@aaronq

OpenStudy (anonymous):

I did it one way, and got 0.982, but then I did it another way, and got 1.57

OpenStudy (anonymous):

Way 1: 4.18 * 103.5 * (32.5−24.5)=c(31.5)(102.4−32.5) 432.63(5)=c(31.5)(69.9) 2163.15=c 2201.85 c=.982

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Idk which way is right??

OpenStudy (anonymous):

@blackops2luvr

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

@RadEn

OpenStudy (anonymous):

2nd way: (103.5) (32.5 - 24.5) (4.181) = (31.5) (102.4 - 32.5) (x) Answer= 1.57

OpenStudy (blackops2luvr):

ummm this is chemistry O_O

OpenStudy (blackops2luvr):

not math

OpenStudy (anonymous):

i know..

OpenStudy (anonymous):

this is the chemistry section.

OpenStudy (blackops2luvr):

ik dat

OpenStudy (anonymous):

do u know anyone on here that knows it?

OpenStudy (anonymous):

@phi

OpenStudy (blackops2luvr):

nop... btw im makeing u a profile pic :)

OpenStudy (anonymous):

hey, i just looked at ur profile pic up close, i know who that dude is, I've seen some pics of him online.. i forgot his name :P

OpenStudy (blackops2luvr):

XD

OpenStudy (blackops2luvr):

der u go :D

OpenStudy (phi):

re-do the first way Way 1: 4.18 * 103.5 * (32.5−24.5)=c(31.5)(102.4−32.5) is correct. But somewhere after this you made a mistake.

OpenStudy (blackops2luvr):

i got a new dog :D

OpenStudy (anonymous):

i want a cartoonish pic like that

OpenStudy (blackops2luvr):

its an american bulldog

OpenStudy (anonymous):

that's so cute, how'd you make that on openstudy

OpenStudy (blackops2luvr):

from a rescue shelter

OpenStudy (anonymous):

thanks Phi

OpenStudy (anonymous):

i like that kind of dog :3

OpenStudy (anonymous):

omg, the thing on the left says: Bump in 167772165: 41 .....

OpenStudy (anonymous):

is that even like real time?? omg

OpenStudy (phi):

here is one problem 32.5−24.5 is not 5

OpenStudy (blackops2luvr):

thats not her.. but shes next to me :D

OpenStudy (blackops2luvr):

her name is phoebe

OpenStudy (anonymous):

is that urs?

OpenStudy (anonymous):

It's big

OpenStudy (blackops2luvr):

.-.

OpenStudy (anonymous):

ooh nice name

OpenStudy (anonymous):

nah im not into the emo theme lol

OpenStudy (blackops2luvr):

XD i drew dat

OpenStudy (anonymous):

I didn't even make the username 'cupcake' , something happened with my account, so openstudy just gave me a random new name

OpenStudy (blackops2luvr):

well a fail drawing

OpenStudy (anonymous):

I gtg do chemistry now :/

OpenStudy (blackops2luvr):

XD ok

OpenStudy (anonymous):

u drew that? awesome ok bye

OpenStudy (blackops2luvr):

buhbye